# Brain-IT: Image Reconstruction from fmri via Brain-Interaction Transformer

### v1

@@ Abstract
Reconstructing images seen by people from their fMRI brain recordings provides a non-invasive window into the human brain. Despite recent progress enabled by diffusion models, current methods often lack faithfulness to the actual seen images. We present “Brain-IT”, a brain-inspired approach that addresses this challenge through a Brain Interaction Transformer (BIT), allowing effective interactions between clusters of functionally-similar brain-voxels. These functional-clusters are shared by all subjects, serving as building blocks for integrating information both within and across brains. All model components are shared by all clusters & subjects, allowing efficient training with a limited amount of data. To guide the image reconstruction, BIT predicts two complementary localized patch-level image features: (i) high-level semantic features which steer the diffusion model toward the correct semantic content of the image; and (ii) low-level structural features which help to initialize the diffusion process with the correct coarse layout of the image. BIT’s design enables direct flow of information from brain-voxel clusters to localized image features. Through these principles, our method achieves image reconstructions from fMRI that faithfully reconstruct the seen images, and surpass current SotA approaches both visually and by standard objective metrics. Moreover, with only 1-hour of fMRI data from a new subject, we achieve results comparable to current methods trained on full 40-hour recordings.
@@ 1 Introduction
Reconstructing visual experiences from brain activity (fMRI-to-image reconstruction) is a key challenge with broad implications for both neuroscience and brain–computer interfaces (Milekovic, 2018; Naci et al., 2012). Such reconstructions may provide a window into visual perception in the brain, enable the study of visual imagery (Cichy et al., 2012; Pearson et al., 2015), reveal dream content (Horikawa et al., 2013; Horikawa & Kamitani, 2017), and even assist in assessing disorders of consciousness (Monti et al., 2010; Owen et al., 2006). In a typical image decoding setting, subjects view natural images while their brain activity is being recorded using functional Magnetic Resonance Imaging (fMRI). This produces paired data of images and their corresponding fMRI scans. The task is then to reconstruct the perceived image from new (test) fMRI signals.
Early work in this domain mapped fMRI signals to handcrafted image features (Kay et al., 2008; Naselaris et al., 2009; Nishimoto et al., 2011), which were then used for image reconstruction. Subsequent studies employed deep learning methods (Beliy et al., 2019; Lin et al., 2019; Shen et al., 2019). Significant progress in reconstruction quality was achieved through the introduction of diffusion models  (Chen et al., 2023; Ozcelik & VanRullen, 2023; Takagi & Nishimoto, 2023). Recent works address the scarcity of fMRI data by integrating information across multiple subjects (Scotti et al., 2024; Gong et al., 2025; Liu et al., 2025), though variability across individuals remains a challenge  (Haxby et al., 2011; Yamada et al., 2015; Wasserman et al., 2024).
@@ 2 Related Work
fMRI-to-image reconstruction is a well-established field, seeing significant progress in the past decade. Early work mapped fMRI to handcrafted image features (Kay et al., 2008; Naselaris et al., 2009; Nishimoto et al., 2011), followed by works that mapped fMRI to deep CNN features (Güçlü & Van Gerven, 2015; Zhang et al., 2018a; Shen et al., 2019). End-to-end methods have emerged (Seeliger et al., 2018; St-Yves & Naselaris, 2018; Beliy et al., 2019), with later ones predicting latent codes of VAEs or GANs (Han et al., 2019; Lin et al., 2019; Mozafari et al., 2020; Qiao et al., 2020; Ren et al., 2021). Recently, diffusion models have improved realism and faithfulness by turning predicted features into high-quality images (Chen et al., 2023; Ozcelik & VanRullen, 2023; Takagi & Nishimoto, 2023). In parallel, the community has increasingly focused on integrating information across multiple subjects to improve generalization under limited data per subject (Scotti et al., 2024; Gong et al., 2025; Huo et al., 2024; Xia et al., 2024b; Ferrante et al., 2024; Liu et al., 2025; Shen et al., 2024). Next, we review related efforts along 3 key axes: predicting image features from fMRI, leveraging cross-subject information, and predicting intermediate low-level images.
@@
Current methods predict image features from fMRI, such as semantic CLIP or VAE latent embeddings. Most use simple linear models or MLPs (Takagi & Nishimoto, 2023; Xia et al., 2024a; Wang et al., 2024b); others build on top with unCLIP diffusion priors (Gong et al., 2025; Scotti et al., 2023). A key limitation of these methods is that they typically compress all voxels into a single global fMRI embedding via a fully connected layer before predicting image features. Since visual information is distributed across multiple distinct yet interconnected brain regions, such a projection fails to fully exploit this distributed nature of the brain. Two recent works introduce spatial voxel groupings (Huo et al., 2024; Shen et al., 2024), which use voxel patches in shared anatomical space. However, they predict a single global image representation, making it difficult to reconstruct localized image information. Our approach addresses these issues by forming functionally shared voxel clusters which are mapped directly to localized image features, avoiding projection to a global fMRI embedding and yielding more accurate image-feature predictions.
@@
Most existing methods regress directly from fMRI to the latent space of the diffusion model (e.g. VAE latent space) using MLPs or linear layers (Scotti et al., 2024; Gong et al., 2025; Ozcelik & VanRullen, 2023). Alternatively, NeuroPictor (Huo et al., 2024) manipulates internal feature maps within the U-Net of Stable Diffusion using features predicted from fMRI. Unlike these approaches, our method introduces a complementary low-level branch that predicts VGG-based representations inspired by LPIPS (Simonyan & Zisserman, 2014; Zhang et al., 2018b) and inverts them through a Deep Image Prior (DIP) framework. This setup results in more faithful low-level predictions, reconstructing global structure and low-level visual features.
@@ 3.1 Image Feature Prediction
The BIT model takes as input the fMRI signals together with the V2C mapping and predicts image features (e.g., adapted CLIP embeddings for the semantic branch, VGG features for the low-level branch). First, BIT transforms brain activations into Brain Tokens, each summarizing the information from the voxels within a voxel cluster. This process incorporates two learnable embeddings to integrate brain information: a per-voxel embedding, which captures each voxel’s functionality, and a per-cluster embedding, which captures each cluster’s overall functionality. This serves as an information-selection bottleneck by determining how voxel information is aggregated into Brain Tokens. The resulting Brain Tokens are then refined via attention layers. Self-attention layers model relationships across Brain Tokens (voxel clusters), while cross-attention layers extract specific image feature information from them. This enables direct information flow from Brain Tokens to predict localized image features.
@@ 3.2 Image Reconstruction
At inference time, the 2 branches are combined: the low-level image initializes the diffusion process, providing coarse image structure, which is then refined into detailed reconstructions with semantic guidance. Reconstructions guided by semantic embeddings alone tend to be semantically consistent but not structurally faithful to the visual stimulus, often failing to preserve low-level details of the seen image such as colors and structure. Kamb & Ganguli (2025) demonstrated that diffusion models generate images in a coarse-to-fine manner: first establishing a global layout and then progressively refining details. Building on this insight, we initialize the diffusion process with a noisy version of the predicted low-level image, which supplies reliable structure. The diffusion model then updates this initialization under semantic conditioning, yielding reconstructions that capture both structure and semantic details, and are more faithful to the seen image. Finally, we apply a refinement stage as in Scotti et al. (2024) in which the image is passed through a pre-trained SDXL diffusion model producing visually enhanced reconstructions (see LABEL:{app:enhance}).
@@
Brain-Diffuser (Ozcelik & VanRullen, 2023) $0.273$ $0.365$ $94.4\%$ $96.6\%$ $91.3\%$ $90.9\%$ $0.728$ $0.421$
Takagi & Nishimoto (2023) $0.246$ $0.410$ $78.9\%$ $85.6\%$ $83.8\%$ $82.1\%$ $0.811$ 0.504
@@ 5 Experimental Results
This section presents the effectiveness of Brain-IT for reconstructing images from brain signals (fMRI). We first briefly describing the experimental framework (Sec. 5.1). We then show quantitative and qualitative results from our full pipeline, showing that it outperforms current SotA methods (Sec. 5.2). Next, we evaluate the transfer learning capabilities of Brain-IT, showing that training on a new subject with only 1 hour of data produces high-quality results, comparable to previous methods trained on the full 40 hours, and significantly better than ones trained on 1 hour (Sec. 5.3).
@@ 6 Ablations and Analysis
This section evaluates the contribution of the two branches (semantic and low-level). Additional ablations of other components and factors are provided in App. A. Those include contribution of: (i) External data, incorporating unlabeled images with no fMRI into training (Sec. A.1); (ii) Functional vs. Anatomical clustering, comparing voxel clusters based on functional representations (as used in Brain-IT) vs. anatomical based clustering (Sec. A.2); (iii) Number of clusters, examining the effect of the number of clusters on reconstruction performance (Sec. A.3).
@@ A.1 Influence of Unlabeled Images
Our method is based on functional clusters, where we assign voxels to clusters using encoding voxel embeddings from the Universal Encoder. An alternative approach is to align all subjects to a common anatomical space and perform local processing of voxels in that space with 2D or 3D patches. For comparison, we construct an anatomical baseline by clustering voxels according to their 3D coordinates in FSaverage space and then apply our pipeline to the resulting clusters. The results below show that our method is flexible and works with different voxel assignments. Within our framework, functional clustering outperforms anatomical clustering.
@@
Anatomical clustering 0.378 0.491 98.1% 99.3% 96.6% 95.6% 0.589 0.336
@@ C.3 Image reconstructions for additional subjects
C.4 Two-Branch Contribution Image Reconstructions

### v2

@@ Abstract
Reconstructing images seen by people from their fMRI brain recordings provides a non-invasive window into the human brain. Despite recent progress enabled by diffusion models, current methods often lack faithfulness to the actual seen images. We present “Brain-IT”, a brain-inspired approach that addresses this challenge through a Brain Interaction Transformer (BIT), allowing effective interactions between clusters of functionally-similar brain-voxels. These functional-clusters are shared by all subjects, serving as building blocks for integrating information both within and across brains. All model components are shared by all clusters & subjects, allowing efficient training with a limited amount of data. To guide the image reconstruction, BIT predicts two complementary localized patch-level image features: (i) high-level semantic features which steer the diffusion model toward the correct semantic content of the image; and (ii) low-level structural features which help to initialize the diffusion process with the correct coarse layout of the image. BIT’s design enables direct flow of information from brain-voxel clusters to localized image features. Through these principles, our method achieves image reconstructions from fMRI that faithfully reconstruct the seen images, and surpass current SotA approaches both visually and by standard objective metrics. Moreover, with only 1-hour of fMRI data from a new subject, we achieve results comparable to current methods trained on full 40-hour recordings. Project page can be found in: https://amitzalcher.github.io/Brain-IT/.
@@ 1 Introduction
Reconstructing visual experiences from brain activity (fMRI-to-image reconstruction) is a key challenge with broad implications for both neuroscience and brain–computer interfaces (Milekovic, 2018; Naci et al., 2012). Such reconstructions may provide a window into visual perception in the brain, enable the study of visual imagery (Cichy et al., 2012; Pearson et al., 2015), reveal dream content (Horikawa et al., 2013; Horikawa and Kamitani, 2017), and even assist in assessing disorders of consciousness (Monti et al., 2010; Owen et al., 2006). In a typical image decoding setting, subjects view natural images while their brain activity is being recorded using functional Magnetic Resonance Imaging (fMRI). This produces paired data of images and their corresponding fMRI scans. The task is then to reconstruct the perceived image from new (test) fMRI signals.
Early work in this domain mapped fMRI signals to handcrafted image features (Kay et al., 2008; Naselaris et al., 2009; Nishimoto et al., 2011), which were then used for image reconstruction. Subsequent studies employed deep learning methods (Beliy et al., 2019; Lin et al., 2019; Shen et al., 2019). Significant progress in reconstruction quality was achieved through the introduction of diffusion models  (Chen et al., 2023; Ozcelik and VanRullen, 2023; Takagi and Nishimoto, 2023). Recent works address the scarcity of fMRI data by integrating information across multiple subjects (Scotti et al., 2024; Gong et al., 2025; Liu et al., 2025), though variability across individuals remains a challenge  (Haxby et al., 2011; Yamada et al., 2015; Wasserman et al., 2024).
@@ 2 Related Work
fMRI-to-image reconstruction is a well-established field, seeing significant progress in the past decade. Early work mapped fMRI to handcrafted image features (Kay et al., 2008; Naselaris et al., 2009; Nishimoto et al., 2011), followed by works that mapped fMRI to deep CNN features (Güçlü and Van Gerven, 2015; Zhang et al., 2018a; Shen et al., 2019). End-to-end methods have emerged (Seeliger et al., 2018; St-Yves and Naselaris, 2018; Beliy et al., 2019), with later ones predicting latent codes of VAEs or GANs (Han et al., 2019; Lin et al., 2019; Mozafari et al., 2020; Qiao et al., 2020; Ren et al., 2021). Recently, diffusion models have improved realism and faithfulness by turning predicted features into high-quality images (Chen et al., 2023; Ozcelik and VanRullen, 2023; Takagi and Nishimoto, 2023). In parallel, the community has increasingly focused on integrating information across multiple subjects to improve generalization under limited data per subject (Scotti et al., 2024; Gong et al., 2025; Huo et al., 2024; Xia et al., 2024b; Ferrante et al., 2024; Liu et al., 2025; Shen et al., 2024). Next, we review related efforts along 3 key axes: predicting image features from fMRI, leveraging cross-subject information, and predicting intermediate low-level images.
@@
Current methods predict image features from fMRI, such as semantic CLIP or VAE latent embeddings. Most use simple linear models or MLPs (Takagi and Nishimoto, 2023; Xia et al., 2024a; Wang et al., 2024b); others build on top with unCLIP diffusion priors (Gong et al., 2025; Scotti et al., 2023). A key limitation of these methods is that they typically compress all voxels into a single global fMRI embedding via a fully connected layer before predicting image features. Since visual information is distributed across multiple distinct yet interconnected brain regions, such a projection fails to fully exploit this distributed nature of the brain. Two recent works introduce spatial voxel groupings (Huo et al., 2024; Shen et al., 2024), which use voxel patches in shared anatomical space. However, they predict a single global image representation, making it difficult to reconstruct localized image information. Our approach addresses these issues by forming functionally shared voxel clusters which are mapped directly to localized image features, avoiding projection to a global fMRI embedding and yielding more accurate image-feature predictions.
@@
Most existing methods regress directly from fMRI to the latent space of the diffusion model (e.g. VAE latent space) using MLPs or linear layers (Scotti et al., 2024; Gong et al., 2025; Ozcelik and VanRullen, 2023). Alternatively, NeuroPictor (Huo et al., 2024) manipulates internal feature maps within the U-Net of Stable Diffusion using features predicted from fMRI. Unlike these approaches, our method introduces a complementary low-level branch that predicts VGG-based representations inspired by LPIPS (Simonyan and Zisserman, 2014; Zhang et al., 2018b) and inverts them through a Deep Image Prior (DIP) framework. This setup results in more faithful low-level predictions, reconstructing global structure and low-level visual features.
@@ 3.1 Image Feature Prediction
The BIT model takes as input the fMRI signals together with the V2C mapping and predicts image features (e.g., adapted CLIP embeddings for the semantic branch, VGG features for the low-level branch). First, BIT transforms brain activations into Brain Tokens, each summarizing the information from the voxels within a voxel cluster. This process incorporates two learnable embeddings to integrate brain information: a per-voxel embedding, which captures each voxel’s functionality, and a per-cluster embedding, which captures each cluster’s overall functionality. This serves as an information-selection bottleneck by determining how voxel information is aggregated into Brain Tokens. The resulting Brain Tokens are then refined via attention layers. Self-attention layers model relationships across Brain Tokens (voxel clusters), while cross-attention layers extract specific image feature information from them. This enables direct information flow from Brain Tokens to predict localized image features as demonstrated in Appendix Sec. C.4, by attention map visualization.
@@ 3.2 Image Reconstruction
At inference time, the 2 branches are combined: the low-level image initializes the diffusion process, providing coarse image structure, which is then refined into detailed reconstructions with semantic guidance. Reconstructions guided by semantic embeddings alone tend to be semantically consistent but not structurally faithful to the visual stimulus, often failing to preserve low-level details of the seen image such as colors and structure. Kamb and Ganguli (2025) demonstrated that diffusion models generate images in a coarse-to-fine manner: first establishing a global layout and then progressively refining details. Building on this insight, we initialize the diffusion process with a noisy version of the predicted low-level image, which supplies reliable structure. The diffusion model then updates this initialization under semantic conditioning, yielding reconstructions that capture both structure and semantic details, and are more faithful to the seen image. Finally, we apply a refinement stage as in Scotti et al. (2024) in which the image is passed through a pre-trained SDXL diffusion model without any text prompt, producing visually enhanced reconstructions (see LABEL:{app:enhance}).
@@
Brain-Diffuser (Ozcelik and VanRullen, 2023) $0.273$ $0.365$ $94.4\%$ $96.6\%$ $91.3\%$ $90.9\%$ $0.728$ $0.421$
Takagi and Nishimoto (2023) $0.246$ $0.410$ $78.9\%$ $85.6\%$ $83.8\%$ $82.1\%$ $0.811$ 0.504
@@ 5 Experimental Results
This section presents the effectiveness of Brain-IT for reconstructing images from brain signals (fMRI). We first briefly describing the experimental framework (Sec. 5.1). We then show quantitative and qualitative results from our full pipeline, showing that it outperforms current SotA methods (Sec. 5.2). Next, we evaluate the transfer learning capabilities of Brain-IT, showing that training on a new subject with only 1 hour of data produces high-quality results, comparable to previous methods trained on the full 40 hours, and significantly better than ones trained on 1 hour (Sec. 5.3). Additionally, we demonstrate the robustness of our model through qualitative evaluation of image reconstructions on “NSD Synthetic” Gifford et al. (2024), which is an out-of-distribution dataset (Fig. S14)
@@ 6 Ablations and Analysis
This section evaluates the contribution of the two branches (semantic and low-level) and provides additional analysis. Further ablations of other components and factors are provided in App. A. Those include contribution of: (i) External data, incorporating unlabeled images with no fMRI into training (Sec. A.1); (ii) Functional vs. Anatomical clustering, comparing voxel clusters based on functional representations (as used in Brain-IT) vs. anatomical based clustering (Sec. A.2); (iii) Number of clusters, examining the effect of the number of clusters on reconstruction performance (Sec. A.3). We also provide an interpretability analysis of Brain-IT via BIT cross-attention maps, showing spatial and semantic selectivity of brain tokens.
@@
Brain-token selectivity.
To better understand how Brain-IT uses fMRI information, we analyze the BIT cross-attention maps between brain tokens (voxel clusters) and query tokens (image-feature positions). Averaging attention across layers and heads reveals that different brain tokens consistently contribute to specific spatial locations in the predicted feature maps, exhibiting a clear contralateral organization. Beyond spatial effects, some tokens preferentially attend to regions associated with particular semantic concepts (e.g., faces, limbs, text), indicating semantic selectivity. Full visualizations are provided in Appendix C.4 (Figs. S12 and S13).
@@ 8 Acknowledgments
I. Ethics statement
We use only publicly available datasets where all the necessary measures to ensure ethical conduct were made. Additionally this work is a continuation of line of works on brain decoding and is a natural extension of previous works.
II. Reproducibility statement
We provide the details of our method in the main paper and additional technical information in App. D. Upon publication, we will release our full code with saved checkpoints to ensure that all results are easily reproducible. We will also share all reconstructed images, enabling researchers to easily compare our method under new evaluation metrics.
@@ A.1 Influence of Unlabeled Images
Our method relies on functional clustering, assigning voxels to clusters based on encoding embeddings from the Universal Encoder. As an anatomical alternative, we aligned all subjects to FSaverage space and clustered voxels using either (i) spatial 3D coordinates or (ii) Schaefer atlas–based cortical parcellations (e.g., Schaefer 7 networks Schaefer et al. (2018)). For the subset of voxels used here, the Schaefer-1000 atlas yielded 198 clusters, and the Schaefer-400 atlas yielded 84 clusters. This demonstrates our framework’s compatibility with multiple clustering strategies, though functional clustering provides superior reconstruction performance.
@@
3D-coordinates clustering 0.378 0.491 98.1% 99.3% 96.6% 95.6% 0.589 0.336
Schaefer-400 parcellation 0.371 0.477 98.3% 99.3% 97.0% 96.1% 0.576 0.333
Schaefer-1000 parcellation 0.378 0.475 98.7% 99.4% 97.4% 96.3% 0.572 0.331
@@ A.5 Two-Branch Contribution
Method Low-Level High-Level
BIT 0.500 0.518 99.4% 99.8% 96.8% 87.5% 0.667 0.462
MLP - InfoNCE loss 0.267 0.481 80.9% 85.0% 77.5% 69.1% 0.897 0.613
MLP - MSE loss 0.182 0.520 54.2% 54.2% 52.0% 51.1% 0.976 0.670
Table T5: Quantitative comparison between BIT and MLP in the low-level branch. We assess the contribution of the Brain-Interaction Transformer (BIT) to low-level feature reconstruction by comparing it against a multilayer perceptron (MLP 1K hidden dim) baseline. The results show that BIT substantially outperforms the MLP predictor in reconstructing low-level visual features. Results are for Subject 1.
@@ C.3 Image reconstructions for additional subjects
C.4 Brain tokens spatial and semantic selectivity
In Brain-IT, each fMRI voxel cluster is represented by a learned brain token, and each predicted image feature is associated with a learned query token. The brain tokens encode the functional activity patterns of specific voxel clusters, while the query tokens index the image features that the model must predict. These two sets of tokens interact through the cross-attention layers of the Brain-Interaction Transformer (BIT): for each query token, BIT attends over all brain tokens and aggregates their information to predict the image feature corresponding to that query.
Each query token is tied to a specific spatial position in an image feature map, so that every predicted feature can be localized in the image. Concretely, for the low-level branch, query tokens correspond to spatial positions in a VGG feature layer (an $A\times A$ grid), and for the semantic branch they correspond to spatial tokens in the CLIP image encoder (e.g., a $16\times 16$ grid of patch tokens). Our goal is to show how different brain tokens contribute to the prediction of different image features via these cross-attention maps, thereby revealing the direct flow of information from functionally defined voxel clusters (brain tokens) to localized image features (query tokens) used for reconstruction. We present two complementary experiments: (i) brain-tokens that contribute to features at specific spatial locations in the image (see Fig. S12); and (ii) brain-tokens that contribute to features capturing specific semantic concepts in the image (see Fig. S13).
Spatial selectivity of brain tokens.
In the first experiment, we visualize the cross-attention map between brain tokens and query tokens by averaging the attention weights across all transformer layers and all attention heads. For each brain token, this yields a 2D attention map over query tokens whose axes correspond to the spatial layout of the image features (an $56\times 56$ grid for a VGG 1_21\_2 layer, and a $16\times 16$ grid of semantic spatial tokens). Each small square in Fig. S12a shows the attention map for one brain token; we arrange these squares to highlight tokens that consistently attend to specific image locations. In Fig. S12b, we assign a unique color to each brain token shown in Fig. S12a, and color all voxels in its corresponding cluster with the same color. This reveals which brain regions most strongly contribute to predicting features at particular image locations. As expected, we observe a clear contralateral organization: brain tokens in the right hemisphere predominantly support features in the left part of the image, and vice versa.
Semantic selectivity of brain tokens.
In the second experiment, we demonstrate semantic selectivity rather than pure spatial position. For each brain token and each image, the corresponding fMRI response induces a different attention map from that token to all query tokens. We again average attention across layers and heads to obtain an attention map per brain token and per image, and then examine, across many images, which spatial regions (and thus which semantic content in those regions) a given brain token most strongly contributes to. We select brain tokens whose attention consistently highlights image regions with a clear semantic meaning, as determined by what typically appears in the attended parts of the images. Fig. S13 illustrates such semantically selective brain tokens, including tokens that preferentially contribute to predicting features for hands, legs, text/words, faces, and other object parts. For each example, we show both the attended image regions and the corresponding voxel cluster on the cortical surface, revealing where in the brain the voxel clusters that support these semantic predictions are located.
Figure S12: Brain Token Spatial Attention Maps (a) Brain tokens’ average attention maps ($6\times 6$ grid, 36 tokens shown), averaged across 1K test images. Each map demonstrates the spatial region in the reconstructed image that the corresponding brain token influences. (b) Brain token brain maps. A color map assigns a color to each brain token according to the position of it’s attention map in the $6\times 6$ grid (a). The brain map shows the voxels corresponding to each brain token, with their color determined by assigned color. This demonstrates the link between a token’s anatomical origin and its attended image location.
Figure S13: Semantic Selectivity of Brain Tokens (a) Semantic attention maps. Each image group showcases the average attention map of a single brain token within the Brain-IT semantic branch, overlaid on the original images. This is evaluated on the test images. We can observe clear semantic selectivity for each brain token (e.g., words, body silhouettes, faces). (b) Brain token brain maps. The brain map shows the anatomical source voxels of the tokens presented in (a), color-coded by the corresponding frame color in (a).
C.5 Reconstructions on NSD-Synthetic
Figure S14: NSD-Synthetic reconstructions results for Subject 1 on the NSD-Synthetic dataset Gifford et al. (2024), an extension of NSD with out-of-distribution stimuli. Brain-IT is applied without any adaptation, demonstrating robust generalization to out-of-distribution stimuli.
C.6 Two-Branch Contribution Image Reconstructions